# Whistle: Data-Efficient Multilingual and Crosslingual Speech Recognition via Weakly Phonetic Supervision

### v1

@@ V Results
In the following, we introduce the experimental results over CV-Lang10, which serves as a common setup for comparing the three MCL-ASR approaches - supervised pre-training with weakly phonetic supervision (Whistle), subword-based supervised pre-training, and wav2vec 2.0 based self-supervised pre-training. The three approaches are described in Section III-A, III-B, and III-C, respectively. An MCL-ASR approach is usually evaluated under two tasks. The first is to recognize utterances from seen languages, i.e., the languages that are included in multilingual pre-training. The second is to recognize utterances from unseen languages, i.e., crosslingual speech recognition, which is often performed by fine-tuning the model obtained from pre-training.
V-A Multilingual pre-training
@@
On the CV-Lang10 dataset, 10 phoneme-based monolingual models are trained, each for a single language and with 90M parameters. Phoneme-based multilingual models (Whistle models) and subword-based multilingual models are trained for comparison. WFST-based decoding are used for all models. The PERs and WERs are shown in Table III and IV respectively. The main observations are as follows.
@@
First, comparing within phoneme-based models, it can be seen that pooling data from multiple languages and training multilingual models clearly reduces PERs over monolingual models, as shown in prior works [14, 31]. Particularly, a single multilingual model (Mult. phoneme L with 543M parameters) performs significantly better than the 10 monolingual separately-trained models (10 * 90M parameters), on averaged PERs over the 10 seen languages. Furthermore, we can see that reductions in WERs can be obtained as well, by phoneme-based multilingual pre-training and WFST-based decoding. Interestingly, in terms of WERs, even the small multilingual model (Mult. phoneme S with 90M parameters) surpasses the monolingual models.
@@
Second, comparing multilingual models based on phonemes and subwords, it is found that the phoneme-based multilingual model (M1) obtains better WERs than the subword-based multilingual model (M4), with close model sizes (around 90M) [^5] The minor difference in model sizes between phoneme-based model and subword-based model (90M vs 92M) is due to the size of the linear layer because of the different alphabet sizes., with 18% WER relative reduction [^6] An exception is that for French, the phoneme-based multilingual model does not outperform the subword-based multilingual model in WER, though the PERs are good. From the statistics of CV-Lang10, we find that the percentage of homophones in the G2P PROLEX of French is the highest (22.5%). The other large percentages of homophones in the 10 langauges in CV-Lang10 is 9.0% for English, 5.2% for Spanish, while others are below 3%. Moreover, it is found that some consonants in French words are usually not pronounced, but they may be pronounced when they are spoken in sentences. The WFST-based decoding with a PROLEX may not be good at capturing these regularities. These issues could be alleviated by developing a better method of decoding from phonemes, which will be explored in future.. This is a fair comparison to answer RQ-1, since both models are trained with the same dataset and the same encoder architecture. Presumably, compared to using subwords which mainly serve for text writing, using phonemes as labels is more natural and better for sound classification, since inherently they are more directly related to describing sounds for languages. Moreover, it can been seen from Figure 2 that data imbalance is more severe in subword supervision than phoneme supervision. From a machine learning perspective, multi-task learning could be severely affected by data imbalance. When data are not well balanced in training, an annoying phenomenon, often observed in subword-based systems, is that high resource languages may suffer from interference and low resource languages may be under-trained, which cause performance degradation [36, 12]. Subword-based systems need special tricks to struggle with data imbalance, such as careful tokenization to appropriately creating the set of tokens [12], human-in-the-loop data mixing in training [36]. In contrast, the superior performances from phoneme-based systems are obtained by training on natural data mixing and adopting the classic IPA symbols that have been matured for describing human sounds for a long time.
@@
Third, we can see clear scaling properties of phoneme-based models - PERs and WERs are consistently reduced for both high-resource and low-resource languages, as the model sizes are increased. Again, remarkably, the performance improvements for different sizes of phoneme-based models are obtained by training on natural data mixing.
TABLE V: PERs and WERs for phoneme-based crosslingual fine-tuning (FT) on Polish. The pre-training dataset is CV-Lang10.
id Model 1h 10h 130h
PER WER PER WER PER WER
O2 Mono. phoneme 86.01 99.98 30.38 13.86 2.82 4.97
M5 W2V (En) phoneme FT 25.76 11.09 16.64  6.75 5.80 4.57
M6 W2V (10 lang) phoneme FT 21.10  7.94 12.65  5.65 6.08 4.44
M7 M1 + phoneme FT 17.96  6.95 10.47  5.27 1.97 4.30
TABLE VI: WERs for subword-based crosslingual fine-tuning (FT) on Polish. The pre-training dataset is CV-Lang10.
id Model 1h 10h 130h
w/o  w w/o  w w/o  w
@@
O3 Mono. subword 98.41 98.38 90.98 59.43 19.38 7.12
M8 W2V (En) subword FT 100 100 45.64 7.08 8.53 3.85
M9 W2V (10 lang) subword FT 99.97 100 36.93 5.71 7.49 3.45
M10 M4 + subword FT 70.13 9.16 31.90 4.89 5.44 3.76
M11 M1 + subword FT 69.50 8.63 31.89 4.83 5.84 3.82
TABLE VII: PERs and WERs for phoneme-based crosslingual fine-tuning (FT) on Indonesian. The pre-training dataset is CV-Lang10.
id Model 1h 10h 20h
PER WER PER WER PER WER
O4 Mono. phoneme 96.52 100 27.30 7.71 5.74 3.28
M12 W2V (En) phoneme FT 31.30 6.73 10.89 3.31 6.84 2.83
M13 W2V (10 lang) phoneme FT 24.91 3.75 10.32 2.79 6.30 2.47
M14 M1 + phoneme FT 21.64 3.27 7.90 2.54 4.79 2.43
TABLE VIII: WERs for subword-based crosslingual fine-tuning (FT) on Indonesian. The pre-training dataset is CV-Lang10.
id Model 1h 10h 20h
w/o  w w/o  w w/o  w
@@
O5 Mono. subword 96.62 96.42 69.57 49.67 31.96 10.85
M15 W2V (En) subword FT 100 100 19.98 5.28 11.68 3.59
M16 W2V (10 lang) subword FT 99.64 99.97 19.08 4.52 12.01 3.15
M17 M4 + subword FT 64.00 23.56 19.41 3.91 13.15 3.07
M18 M1 + subword FT 67.71 24.57 18.21 3.59 12.48 2.92
V-B Crosslingual fine-tuning
Over the CV-Lang10 dataset, we obtain the phoneme-based supervised pre-trained model (M1), which can be further fine-tuned with either phoneme labels or subword labels for crosslingual speech recognition. The subword-based supervised pre-trained model (M4) is fine-tuned with subword labels for crosslingual speech recognition. The wav2vec 2.0 models, “W2V (10 lang)” and “W2V (En)”, can be fine-tuned with either phoneme labels or subword labels for crosslingual speech recognition. The four pre-trained models used in the crosslingual experiments all have the same model size (around 90M parameters). On the four pre-trained models, we perform full-parameter fine-tuning, except that for the two wav2vec 2.0 based pre-trained models, the convolutional feature encoder are frozen.
To test different multilingual pre-trained models for crosslingual speech recognition, we conduct phoneme-based and subword-based crosslingual fine-tuning on unseen languages. The training data from an unseen language is divided into three scales to simulate different resource scenarios, while the test and validation data remain unchanged.
The first unseen language is Polish. Polish has 31 phonemes contained in CV-Lang10 and 4 unseen phonemes. The training data is divided into three scales: 1 hour, 10 hours, and full (130 hours). Combining Table V and Table VI, we have the following main observations.
  • In the low-resource scenario with 1-hour Polish training data, phoneme pre-training (PT) followed by phoneme fine-tuning (FT) performs the best (6.95). Results with phoneme PT are much better than those with subword PT, which clearly shows the advantage of phonetic supervision in representation learning from multilingual data (RQ-1). When comparing phoneme PT and wav2vec 2.0 PT (M7 vs M6, M11 vs M9), phoneme PT shows obvious superiority (RQ-2).
  • With the full Polish training data, the wav2vec 2.0 PT models start to perform well, surpassing results with both subword PT and phoneme PT (3.45 ¡ 3.76 ¡ 3.82). This may reflect some benefit of wav2vec 2.0 PT when fine-tuned with abundant labels, but such top-performing result with wav2vec 2.0 PT is not observed in Indonesian experiments, as shown below.
The second unseen language is Indonesian. All 35 phonemes of Indonesian are contained in CV-Lang10. But Indonesian belongs to the Austronesian language family, which are somewhat more different from CV-Lang10, and only 20 hours of training data are available. These make crosslingual fine-tuning for Indonesian more challenging. The training data is divided into three scales: 1 hour, 10 hours, and full (20 hours).
From Table VII and Table VIII for Indonesian, the observations are similar to those for Polish. In the more challenging scenario with larger linguistic difference and less training data, the advantages of phoneme PT followed by phoneme FT are more obvious, across all the three scales of data settings. It seems that when training data are more limited, the better results can be obtained by phoneme supervision, compared to subword supervision and self-supervision. When the amount of crosslingual training data increases, the performance gaps between phoneme supervision, subword supervision and self-supervision may diminish. Presumably, the fine-tuning with abundant data behaves like end-to-end monolingual training and the effect of different PT methods may become weak.
@@ VI-B Test of catastrophic forgetting
In previous sections, we show the advantage of multilingual pre-trained models by phoneme supervision over those by subword supervision for recognizing seen and unseen languages. We see that after a pre-trained multilingual model is fine-tuned over data from a new language, the fine-tuned multilingual model can recognize speech from the new language. Then, to what degree the performance of the fine-tuned multilingual model on previous seen languages would be affected? This is an interesting question for continual pre-training of multilingual models to support more new languages, a question related to catastrophic forgetting of neural network based models [58]. A complete investigation into continual pre-training of multilingual models is outside the scope of this paper. Here we present a preliminary examination of the two approaches, phoneme or subword-based multilingual models, in overcoming catastrophic forgetting.
@@
TABLE IX: Test of catastrophic forgetting for the multilingual models, pre-trained over CV-Lang10 and fine-tuned on 10 minutes of a new language (Polish). WARD denotes word accuracy relative degradation of the averaged WER over the ten old languages in CV-Lang10.
id Model pl en fr es it ru nl tr ky sv tt Avg. WARD
M19 M1 + 10min phoneme FT 11.0 68.5 69.3 57.1 50.3 48.3 60.9 31.8 58.4 42.3 33.0 52.0 48
M20 M4 + 10min subword FT 93.2 92.2 95.0 92.5 92.5 262.5 103.6 241.5 125.9 180.5 254.4 154.1 160
@@ VI-C Training efficiency
TABLE X: Training efficiency of phoneme-based and subword-based pre-training (PT) and fine-tuning (FT).
id Model Batch size Epochs for converging
M1 phoneme PT 640 63
M11 M1 + pl subword FT 320 195
M4 subword PT 640 83
M10 M4 + pl subword FT 320 223

### v2

@@ V Results
In the following, we introduce the experimental results over CV-Lang10, which serves as a common setup for comparing the three MCL-ASR approaches - supervised pretraining with weakly phonetic supervision (Whistle), subword-based supervised pretraining, and wav2vec 2.0 based self-supervised pretraining. The three approaches are described in Section III-A, III-B, and III-C, respectively. An MCL-ASR approach is usually evaluated under two tasks. The first is to recognize utterances from seen languages, i.e., the languages that are included in multilingual pretraining. The second is to recognize utterances from unseen languages, i.e., crosslingual speech recognition, which is often performed by finetuning the model obtained from pretraining.
V-A Multilingual pretraining
@@
On the CV-Lang10 dataset, 10 phoneme-based monolingual models are trained, each for a single language and with 90M parameters. Phoneme-based multilingual models (Whistle models) and subword-based multilingual models are trained for comparison. WFST-based decoding are used for all models. The PERs and WERs are shown in Table IV. The main observations are as follows.
@@
1) Comparing within phoneme-based models, it can be seen that pooling data from multiple languages and training multilingual models clearly reduces PERs over monolingual models, as shown in prior works [14, 31]. Particularly, a single multilingual model (Mult. phoneme L with 543M parameters) performs significantly better than the 10 monolingual separately-trained models (10 * 90M parameters), on averaged PERs over the 10 seen languages. Furthermore, we can see that reductions in WERs can be obtained as well, by phoneme-based multilingual pretraining and WFST-based decoding. Interestingly, in terms of WERs, even the small multilingual model (Mult. phoneme S with 90M parameters) surpasses the monolingual models.
@@
2) Comparing the phoneme-based and subword-based multilingual models, it is found that the phoneme-based multilingual model (M1) obtains better WERs than the subword-based multilingual model (M4), with a relative WER reduction of 18% [^7] An exception is that for French, the phoneme-based multilingual model does not outperform the subword-based multilingual model in WER, though the PERs are good. From the statistics of CV-Lang10, we find that the percentage of homophones in the G2P PROLEX of French is the highest (22.5%). The other large percentages of homophones in the 10 langauges in CV-Lang10 is 9.0% for English, 5.2% for Spanish, while others are below 3%. Moreover, it is found that some consonants in French words are usually not pronounced, but they may be pronounced when they are spoken in sentences. The WFST-based decoding with a PROLEX may not be good at capturing these regularities. These issues could be alleviated by developing a better method of decoding from phonemes, which will be explored in future.. Both models are trained with the same dataset and the same encoder architecture, with close model sizes (around 90M) [^8] The minor difference in model sizes between phoneme-based model and subword-based model (90M vs 92M) is due to the size of the linear layer because of the different alphabet sizes.. This is a fair comparison to answer RQ-1, representing better multilingual data-efficiency of phonetic supervision over graphemic supervision. Intuitively, compared to using subwords which mainly serve for text writing, using phonemes as labels is more natural and better for sound classification, since inherently they are more directly related to describing sounds for languages.
In the following, we provide two points to understand why phonetic supervision obtains better multilingual data-efficiency than graphemic supervision, from the perspectives of data balance and data augmentation respectively. First, graphemic supervision suffers from a more severe data imbalance than phoneme supervision, which can be clearly seen from Figure 2. From a machine learning perspective, multi-task learning could be severely affected by data imbalance. When data are not well balanced in training, an annoying phenomenon, often observed in subword-based systems, is that high resource languages may suffer from interference and low resource languages may be under-trained, which cause performance degradation [36, 12]. Subword-based systems need special tricks to struggle with data imbalance, such as careful tokenization to appropriately creating the set of tokens [12], human-in-the-loop data mixing in training [36]. In contrast, the superior performances from phoneme-based systems are obtained by training on natural data mixing and adopting the classic IPA symbols that have been matured for describing human sounds for a long time.
Figure 3: Relative reduction in WER (RRWER) (comparing phoneme pretraining (M1) against subword pretraining (M4) in multilingual speech recognition), as a function of relative increase in phoneme occurrences (RIPO), for the ten languages in CV-Lang10. The figure shows the line of best linear fit: $\text{RRWER}=0.39\times\text{RIPO}+6.6$.
Second, phonetic supervision enables a more efficient data sharing than graphemic supervision in multilingual training. Given a certain amount of training data from a language in multilingual training, the training data for a phoneme in this language is actually augmented in multilingual training, if this phoneme is also occurred in other languages. The more sharing in other languages, the more training data is augmented. This effect can be viewed as an implicit kind of data augmentation. Considering that phonemes are more shared between different languages than subwords, such data augmentation is stronger in phoneme-based multilingual pretraining than in subword-based multilingual pretraining. Presumably, this explains the better performance of phonetic supervision than graphemic supervision, when both are trained over the same amount of data, i.e. better multilingual data-efficiency. We can perform a numerical analysis. For each language, we calculate a base size, i.e., the sum of the number of occurrences of the phonemes in this language in its own data, and an augmented size, i.e., the sum of the number of occurrences of these phonemes by counting over the entire multilingual data. We then calculate the relative increase in phoneme occurrences (RIPO), by comparing the augmented size against the base size. Meanwhile, for each seen language, we calculate the relative reduction in WER (RRWER), when phoneme pretraining (M1) is compared against subword pretraining (M4). Figure 3 shows RRWER as a function of RIPO for the ten languages in CV-Lang10. We can see that RRWER is positively correlated with RIPO. Low-resource languages obtain more data sharing from multilingual phoneme-based pretraining, enabling them to achieve more relative WER reduction compared to subword-based pretraining.
@@
3) We can see clear scaling properties of phoneme-based models - PERs and WERs are consistently reduced for both high-resource and low-resource languages, as the model sizes are increased. Again, remarkably, the performance improvements for different sizes of phoneme-based models are obtained by training on natural data mixing.
V-B Crosslingual finetuning
Over the CV-Lang10 dataset, we obtain the phoneme-based supervised pretrained model (M1), which can be further finetuned with either phoneme labels or subword labels for crosslingual speech recognition. The subword-based supervised pretrained model (M4) is finetuned with subword labels for crosslingual speech recognition. The wav2vec 2.0 models, “Wav2vec2 (10 lang)” and “Wav2vec2 (En)”, can be finetuned with either phoneme labels or subword labels for crosslingual speech recognition. The four pretrained models used in the crosslingual experiments all have the same model size (around 90M parameters). On the four pretrained models, we perform full-parameter finetuning, except that for the two wav2vec 2.0 based pretrained models, the convolutional feature encoder are frozen.
To test different multilingual pretrained models for crosslingual speech recognition, we conduct phoneme-based and subword-based crosslingual finetuning on unseen languages. The training data from an unseen language is divided into three scales to simulate different resource scenarios, while the test and validation data remain unchanged.
The first unseen language is Polish. Polish has 31 phonemes contained in CV-Lang10 and 4 unseen phonemes. The training data is divided into three scales: 1 hour, 10 hours, and full (130 hours). From Table V, we have the following main observations.
  • In the low-resource scenario with 1-hour Polish training data, phoneme pretraining (PT) followed by phoneme finetuning (FT) performs the best (6.95). Results with phoneme PT are much better than those with subword PT, which clearly shows the advantage of phonetic supervision in representation learning from multilingual data (RQ-1). When comparing phoneme PT and wav2vec 2.0 PT (M8 vs M6), phoneme PT shows obvious superiority (RQ-2).
  • With the full Polish training data, the wav2vec 2.0 PT models start to perform well, surpassing results with both subword PT and phoneme PT (3.45 ¡ 3.76 ¡ 3.82). This may reflect some benefit of wav2vec 2.0 PT when finetuned with abundant labels, but such top-performing result with wav2vec 2.0 PT is not observed in Indonesian experiments, as shown below.
The second unseen language is Indonesian. All 35 phonemes of Indonesian are contained in CV-Lang10. But Indonesian belongs to the Austronesian language family, which are somewhat more different from CV-Lang10, and only 20 hours of training data are available. These make crosslingual finetuning for Indonesian more challenging. The training data is divided into three scales: 1 hour, 10 hours, and full (20 hours).
From Table VI for Indonesian, the observations are similar to those for Polish. In the more challenging scenario with larger linguistic difference and less training data, the advantages of phoneme PT followed by phoneme FT are more obvious, across all the three scales of data settings. It seems that when training data are more limited, the better results can be obtained by phoneme supervision, compared to subword supervision and self-supervision. When the amount of crosslingual training data increases, the performance gaps between phoneme supervision, subword supervision and self-supervision may diminish. Presumably, the finetuning with abundant data behaves like end-to-end monolingual training and the effect of different PT methods may become weak. In summary, we find that compared to both graphemic supervision and self-supervision, phonetic supervision excels in crosslingual data-efficiency. The efficiency advantage is more significant when the data are more limited and may diminish when the finetuning data are abundant.
TABLE V: Error rates for phoneme-based and subword-based crosslingual finetuning (FT) on Polish. The pretraining (PT) dataset is CV-Lang10. For phoneme FT, we report PERs for beam search without PROLEX and WERs for WFST decoding, respectively. For subword FT, we report WERs for beam search without LM and WFST decoding with LM, respectively.
ID Pretrained Model Phoneme FT (PER/WER) Subword FT (WER)
1 hour 10 hour 130 hour 1 hour 10 hour 130 hour
PER WER PER WER PER WER w/o  w w/o  w w/o  w
@@ V-B Crosslingual finetuning
O2 Mono. 86.01 99.98 30.38 13.86 2.82 4.97 98.41 98.38 90.98 59.43 19.38 7.12
M5 Wav2vec2 (En) 25.76 11.09 16.64  6.75 5.80 4.57 100 100 45.64 7.08 8.53 3.85
M6 Wav2vec2 (10 lang) 21.10  7.94 12.65  5.65 6.08 4.44 99.97 100 36.93 5.71 7.49 3.45
M7 M4 (subword PT) - - - - - - 70.13 9.16 31.90 4.89 5.44 3.76
M8 M1 (phoneme PT) 17.96  6.95 10.47  5.27 1.97 4.30 69.50 8.63 31.89 4.83 5.84 3.82
TABLE VI: Error rates for phoneme-based and subword-based crosslingual finetuning (FT) on Indonesian. The pretraining (PT) dataset is CV-Lang10.
ID Pretrained Model Phoneme FT (PER/WER) Subword FT (WER)
1 hour 10 hour 20 hour 1 hour 10 hour 20 hour
PER WER PER WER PER WER w/o  w w/o  w w/o  w
@@
O3 Mono. 96.52 100 27.30 7.71 5.74 3.28 96.62 96.42 69.57 49.67 31.96 10.85
M9 Wav2vec2 (En) 31.30 6.73 10.89 3.31 6.84 2.83 100 100 19.98 5.28 11.68 3.59
M10 Wav2vec2 (10 lang) 24.91 3.75 10.32 2.79 6.30 2.47 99.64 99.97 19.08 4.52 12.01 3.15
M11 M4 (subword PT) - - - - - - 64.00 23.56 19.41 3.91 13.15 3.07
M12 M1 (phoneme PT) 21.64 3.27 7.90 2.54 4.79 2.43 67.71 24.57 18.21 3.59 12.48 2.92
@@ VI-A Analysis of embeddings
TABLE VII: Test of catastrophic forgetting for the multilingual models, pretrained over CV-Lang10 and finetuned on 10 minutes of a new language (Polish). WARD denotes word accuracy relative degradation of the averaged WER over the ten old languages in CV-Lang10.
Model pl en fr es it ru nl tr ky sv tt Avg. WARD
M1 + 10min phoneme FT 11.0 68.5 69.3 57.1 50.3 48.3 60.9 31.8 58.4 42.3 33.0 52.0 48
M4 + 10min subword FT 93.2 92.2 95.0 92.5 92.5 262.5 103.6 241.5 125.9 180.5 254.4 154.1 160
TABLE VIII: Training efficiency of phoneme-based and subword-based pretraining (PT) and finetuning (FT).
Model Batch size Epochs for converging
M1 640 63
M1 + pl subword FT 320 195
M4 640 83
M4 + pl subword FT 320 223
@@ VI-B Test of catastrophic forgetting
In previous sections, we show the advantage of multilingual pretrained models by phoneme supervision over those by subword supervision for recognizing seen and unseen languages. We see that after a pretrained multilingual model is finetuned over data from a new language, the finetuned multilingual model can recognize speech from the new language. Then, to what degree the performance of the finetuned multilingual model on previous seen languages would be affected? This is an interesting question for continual pretraining of multilingual models to support more new languages, a question related to catastrophic forgetting of neural network based models [63]. A complete investigation into continual pretraining of multilingual models is outside the scope of this paper. Here we present a preliminary examination of the two approaches, phoneme or subword-based multilingual models, in overcoming catastrophic forgetting.
@@ VI-C Training efficiency
VI-D Last linear layer initialization in finetuning from subword-based pretrained models
As introduced in Section III, the row vectors from the weight matrix of the last linear layer can be viewed as embedding vectors for output units, which can be phonemes or subwords. Denote the union of the unit inventories from the seen languages in pretraining by $V_{\text{multi}}$, and the unit inventory for an unseen language in finetuning by $V_{\text{cross}}$, respectively. In crosslingual finetuning, the parameters corresponding to seen units in $V_{\text{multi}}\cap V_{\text{cross}}$ can be initialized by directly copying from the pretrained model and those parameters for unseen units are randomly initialized. This is exactly what we do in finetuning from phoneme-based pretrained models in our experiments.
For finetuning from subword-based pretrained models, this similar initialization scheme can be used as well, i.e., copying for seen subwords and random initialization for the remaining subwords. As suggested by a referee, we compare this scheme to the completely random initialization scheme. The comparison results are shown in Table IX. It can be seen that the initialization scheme with copying for seen units [^10] Among the 500 BPE subwords for Polish, there are 202 seen subwords in the pretraining data (i.e. CV-Lang10); and for the 500 BPE subwords for Indonesian, there are 234 seen subwords in CV-Lang10. performs worse than random initialization for finetuning from subword-based pretrained models. Presumably, this is because when the same subword occurs in different languages, the sounds pronounced can often vary significantly [^11] For example, the subword “nat” in the English word “nature” is pronounced as [net], while the subword “nat” in the Polish word “natomiast” is pronounced as [nat].. The model may be confused in learning. Initialization with copying from seen units does not improve the learning in this case. Therefore, random initialization is the common practice in finetuning from subword-based pretrained models, which is exactly what we do in our experiments and reflects its best performance. The comparison of finetuning from phoneme-based and subword based pretrained models in our experiments as shown in Table V and Table VI is sound. This further strengthens our basic observation: the essential role of subwords is primarily for the writing of a language, rather than for describing and distinguishing sounds, for which phonemes are defined.
TABLE IX: Word error rates for crosslingual finetuning from the subword-based pretrained model (M4) with different initialization schemes for the Last linear layer. Note that the results by random initialized are taken from Table V and Table VI, which are shown here for clear side-by-side comparison.
Exp Polish Indonesian
1 hour 10 hour 130 hour 1 hour 10 hour 20 hour
w/o  w w/o  w w/o  w w/o  w w/o  w w/o  w
LM LM LM LM LM LM LM LM LM LM LM LM
Random initialization 70.13 9.16 31.90 4.89 5.44 3.76 64.00 23.56 19.41 3.91 13.15 3.07
Copying for seen units 71.39 12.73 32.11 5.39 6.54 4.17 57.08 24.84 19.72 4.99 13.40 3.40